### v1

@@ 1 Introduction
The naive Monte Carlo estimator for KL divergence (Eq. 1) involves sampling $M$ strings $\boldsymbol{Y}^{(1)},\ldots,\boldsymbol{Y}^{(M)}\mathrel{\stackrel{{\scriptstyle\textnormal{i.i.d.}}}{{\sim}}}p$ and then averaging $\log\frac{p(\boldsymbol{Y}^{(m)})}{q(\boldsymbol{Y}^{(m)})}$. Even though this estimator is unbiased, it often exhibits high variance. More pathologically, the naive Monte Carlo estimator can result in negative estimates of KL, which may be undesirable in practice. To address these issues, practitioners adopt alternative techniques to ensure non-negativity. For example, Schulman (2020) proposes an unbiased, non-negative KL estimator that is widely used in practice (von Werra et al., 2020; Havrilla et al., 2023). However, it offers no theoretical variance improvement over the Monte Carlo estimator and, as we show empirically, can exhibit unbounded variance.
@@ 2.1 Language Modeling
$\vec{p}(\boldsymbol{x})\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\sum_{\boldsymbol{y}\in\Sigma^{*}}p(\boldsymbol{x}\boldsymbol{y}),$ (2)
@@
$\vec{p}(\boldsymbol{x}\mid\boldsymbol{y})=\frac{\vec{p}(\boldsymbol{x}\boldsymbol{y})}{\vec{p}(\boldsymbol{y})}$ (3)
where we additionally define $\vec{p}(\textsc{eos}\mid\boldsymbol{y})\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\frac{p(\boldsymbol{y})}{\vec{p}(\boldsymbol{y})}$.
@@
where $\boldsymbol{y}_{<n}\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}y_{1}\cdots y_{n-1}$ and $\textsc{eos}\notin\Sigma$ is a distinguished end-of-string symbol. Let $\overline{\Sigma}\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\Sigma\cup\{\textsc{eos}\}$. In Eq. 4, each $\vec{p}(\cdot\mid\boldsymbol{y}_{<n})$ can fruitfully be viewed as a distribution over $\overline{\Sigma}$. Despite the overloading of the notation, whether $\vec{p}(\cdot\mid\boldsymbol{y}_{<n})$ refers to a prefix probability, a function which takes an argument from $\Sigma^{*}$, or a distribution over $\overline{\Sigma}$ will always be clear from context. Throughout this paper, we use $\boldsymbol{Y}$ to represent the string-valued random variable sampled from $p$. When taking $M$ samples from $p$, we use $\boldsymbol{Y}^{(m)}$ to denote the $m^{\text{th}}$ sample.
@@
$\textsc{kl}(p\mid\mid q)\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\sum_{\boldsymbol{y}\in\Sigma^{*}}p(\boldsymbol{y})\log\frac{p(\boldsymbol{y})}{q(\boldsymbol{y})}=\operatorname*{\mathbb{E}}\left[f(\boldsymbol{Y})\right],$ (5)
where the function that we want to estimate is denoted with $f$, which in this case is $\log\frac{p(\boldsymbol{Y})}{q(\boldsymbol{Y})}$. Indeed, for the remainder of the paper, we define $f(\boldsymbol{Y})\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\log\frac{p(\boldsymbol{Y})}{q(\boldsymbol{Y})}$. Note that while the exact KL value is always non-negative, $f(\boldsymbol{Y})$ may be positive or negative. Importantly, the exact computation of Eq. 5 involves iterating over all strings in $\Sigma^{*}$.
@@ 2.2 Monte Carlo Estimation of KL
$=\frac{1}{M}\sum_{m=1}^{M}f(\boldsymbol{Y}^{(m)}),$ (6b)
where $\boldsymbol{Y}^{(1)},\ldots,\boldsymbol{Y}^{(M)}\mathrel{\stackrel{{\scriptstyle\textnormal{i.i.d.}}}{{\sim}}}p$. Throughout the paper, we assume that the KL divergence is finite, i.e., $\textsc{kl}(p\mid\mid q)<\infty$. It is straightforward to show $\delta_{\textsc{mc}}$ is unbiased, i.e., $\operatorname*{\mathbb{E}}[\delta_{\textsc{mc}}]=\textsc{kl}(p\mid\mid q)$. The variance of this estimator is
$\mathop{\mathrm{Var}}\left[{\normalcolor\delta_{\textsc{mc}}}\right]$ $=\frac{1}{M^{2}}\mathop{\mathrm{Var}}\left[{\normalcolor\sum_{m=1}^{M}f(\boldsymbol{Y}^{(m)})}\right]$ (7a)
$=\frac{1}{M}\mathop{\mathrm{Var}}\left[{\normalcolor f(\boldsymbol{Y})}\right].$ (7b)
@@ 2.3 Control Variate Monte Carlo
A general technique to reduce the variance of any estimator is by introducing control variates (§8.2; Ross, 2002). In the context of estimating the KL divergence between two language models, control variate Monte Carlo was popularized by Schulman (2020) and is commonly implemented in many RLHF libraries (Havrilla et al., 2023; Hu et al., 2024; Sheng et al., 2024). Formally, a control variate is any function $g\colon\Sigma^{*}\to\mathbb{R}$ for which $G\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\operatorname*{\mathbb{E}}\left[g(\boldsymbol{Y})\right]$ can be efficiently computed. [^4] One could also consider control variates of the form $g\colon\Sigma^{*}\to\mathbb{R}^{d}$ for $d>1$ (Geffner and Domke, 2018). Armed with a control variate, we define the control variate Monte Carlo as []
@@
which directly translates to reducing the variance of the Monte Carlo estimator (compare Eq. 7b and Eq. 10c). The greater the correlation between $f$ and $g$, the more we reduce the variance. Note that the value of $\alpha$ may be estimated from a pilot sample when it cannot be computed analytically. [^5] Note that the control variate method can be straightforwardly extended to support multiple control variates.
@@ 3 Rao–Blackwellized Monte Carlo
$\delta_{\textsc{mc}}^{N}$ =$\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\frac{1}{M}\sum_{m=1}^{M}\log\frac{\vec{p}(\overline{\boldsymbol{Y}}_{\leq N}^{(m)})}{\vec{q}(\overline{\boldsymbol{Y}}_{\leq N}^{(m)})}$ (15a)
@@
$=\sum_{n=1}^{N}\underbrace{\frac{1}{M}\sum_{m=1}^{M}\log\frac{\vec{p}(\overline{Y}_{n}^{(m)}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})}{\vec{q}(\overline{Y}_{n}^{(m)}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})}}_{\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\delta_{\textsc{mc}}^{n}}$ (15d)
@@
$\delta_{\textsc{rb}}^{N}$ =$\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\sum_{n=1}^{N}\operatorname*{\mathbb{E}}\left[\delta_{\textsc{mc}}^{n}\mid T_{n}\right]$ (17a)
$=\sum_{n=1}^{N}\frac{1}{M}\sum_{m=1}^{M}\operatorname*{\mathbb{E}}_{\overline{Y}_{n}}\left[\log\frac{\vec{p}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})}{\vec{q}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})}\right],$ (17b)
where the expectation in Eq. 17b is the exact KL divergence between the conditional distributions and is computed as: $\sum_{\overline{Y}_{n}\in\overline{\Sigma}}\vec{p}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})\log\frac{\vec{p}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})}{\vec{q}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})}$. The following theorem establishes that $\delta_{\textsc{rb}}^{N}$ is an unbiased estimator of $\delta_{\textsc{mc}}^{N}$, and furthermore, that $\delta_{\textsc{rb}}^{N}$ has lower or equal variance compared to $\delta_{\textsc{mc}}^{N}$.
@@
With regard to the estimator $\delta_{\textsc{rb}}^{N}$ defined in Eq. 17a, the following two properties hold
@@
We then define the general case as $\delta_{\textsc{rb}}\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\lim_{N\rightarrow\infty}\delta_{\textsc{rb}}^{N}$. The following theorem shows that $\delta_{\textsc{rb}}$ is an unbiased estimator of the KL divergence with variance less than or equal to $\delta_{\textsc{mc}}$.
@@
Notably, $\delta_{\textsc{rb}}$ is guaranteed to be non-negative. Each step-wise estimator computes the exact KL divergence between the two distributions, conditioned on the sampled context $\boldsymbol{y}_{<n}^{(m)}$. Since all terms in Eq. 17a are non-negative, $\delta_{\textsc{rb}}$ remains non-negative as well.
@@
At first glance, computing the Rao–Blackwellized estimator in Eq. 17a might seem more computationally expensive than the standard Monte Carlo estimator in Eq. 6a, since it requires evaluating the exact KL divergence at each token position. However, this added computation does not change the overall runtime complexity. To sample $M$ strings of at most length $N$, we perform a forward pass through the language model for $N$ tokens, followed by the final transformation and a softmax. Assuming that the model representations live in $\mathbb{R}^{d}$, the next token prediction step requires ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathcal{O}\!\left(MNd|\Sigma|\right)}$ operations, where $|\Sigma|$ is the size of the alphabet. Thus, the total time complexity of sampling $M$ strings is ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathcal{O}\!\left(MT_{N}+MNd|\Sigma|\right)}$, where $T_{N}$ denotes the time complexity of performing a forward pass for a single string of length $N$. For KL estimation, computing the MC estimator requires an additional ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathcal{O}\!\left(MN\right)}$ operations, while the Rao–Blackwellized estimator requires ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathcal{O}\!\left(MN|\Sigma|\right)}$. Therefore, despite the additional computation in the Rao–Blackwellized estimator, its overall runtime complexity is the same as the standard Monte Carlo estimator. In general, $N$ is itself a random variable. Thus, a more fine-grained, randomized analyze would yield the bound ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathcal{O}\!\left(MT_{\mathbb{E}[N]}+M\mathbb{E}[N]d|\Sigma|\right)}$ where $\mathbb{E}[N]$ is the expected string length under $p$.
@@ 4 Estimating the Gradient
$\nabla_{\boldsymbol{\theta}}\textsc{kl}$ $=\sum_{\boldsymbol{y}\in\Sigma^{*}}\nabla_{\boldsymbol{\theta}}p_{\boldsymbol{\theta}}(\boldsymbol{y})\log\frac{p_{\boldsymbol{\theta}}(\boldsymbol{y})}{q(\boldsymbol{y})}+\nabla_{\boldsymbol{\theta}}p_{\boldsymbol{\theta}}(\boldsymbol{y})$ (19a)
$=\operatorname*{\mathbb{E}}\left[\left(\log\frac{p_{\boldsymbol{\theta}}(\boldsymbol{Y})}{q(\boldsymbol{Y})}+1\right)\nabla_{\boldsymbol{\theta}}\log p_{\boldsymbol{\theta}}(\boldsymbol{Y})\right].$ (19b)
@@
$\begin{split}\nabla_{\boldsymbol{\theta}}\textsc{kl}(p_{\boldsymbol{\theta}}\mid\mid q)=\sum_{\boldsymbol{y}\in\Sigma^{*}}\vec{p}_{\boldsymbol{\theta}}(\boldsymbol{y})\operatorname*{\mathbb{E}}_{\overline{Y}}\Big{[}\log\frac{\vec{p}_{\boldsymbol{\theta}}(\overline{Y}\mid\boldsymbol{y})}{\vec{q}(\overline{Y}\mid\boldsymbol{y})}\\ \cdot\left(\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\overline{Y}\mid\boldsymbol{y})+\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\boldsymbol{y})\right)\Big{]}.\end{split}$ (23)
@@
$\begin{split}\nabla\delta_{\textsc{rb}}=\frac{1}{M}\sum_{m=1}^{M}\lim_{N\rightarrow\infty}\sum_{n=1}^{N}\operatorname*{\mathbb{E}}_{\overline{Y}_{n}}\Big{[}\log\frac{\vec{p}_{\boldsymbol{\theta}}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})}{\vec{q}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})}\\ \cdot\left(\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\overline{\boldsymbol{Y}}^{(m)}_{<n})+\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})\right)\Big{]}.\end{split}$ (24)
@@
Let $p_{\boldsymbol{\theta}}$ and $q$ be two language models over $\Sigma$ and $\vec{p}_{\boldsymbol{\theta}}$ the prefix probability function of $p_{\boldsymbol{\theta}}$. Let $\nabla\delta_{\textsc{mc}}$ be the Monte Carlo estimator of the true gradient, denoted with $G\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\nabla_{\boldsymbol{\theta}}\textsc{kl}(p_{\boldsymbol{\theta}}\mid\mid q)$, as defined in Eq. 21 and $\nabla\delta_{\textsc{rb}}$ be the Rao–Blackwellized estimator of the gradient defined in Eq. 24. Assuming $\mathrm{H}(p_{\boldsymbol{\theta}},q)<\infty,\mathop{\mathrm{Var}}\left[{\normalcolor\nabla\delta_{\textsc{rb}}}\right]<\infty,\mathop{\mathrm{Var}}\left[{\normalcolor\nabla\delta_{\textsc{mc}}}\right]<\infty$, we have

### v2

@@ 1 Introduction
The naive Monte Carlo estimator for KL divergence (Eq. 1) involves sampling $M$ strings $\boldsymbol{Y}^{(1)},\ldots,\boldsymbol{Y}^{(M)}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny i.i.d.}}}{{\sim}}}p$ and then averaging $\log\frac{p(\boldsymbol{Y}^{(m)})}{q(\boldsymbol{Y}^{(m)})}$. Even though this estimator is unbiased, it often exhibits high variance. More pathologically, the naive Monte Carlo estimator can result in negative estimates of KL, which may be undesirable in practice. To address these issues, practitioners adopt alternative techniques to ensure non-negativity. For example, Schulman (2020) proposes an unbiased, non-negative KL estimator that is widely used in practice (von Werra et al., 2020; Havrilla et al., 2023). However, it offers no theoretical variance improvement over the Monte Carlo estimator and, as we show empirically, can exhibit unbounded variance.
@@ 2.1 Language Modeling
$\vec{p}(\boldsymbol{x})\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\sum_{\boldsymbol{y}\in\Sigma^{*}}p(\boldsymbol{x}\boldsymbol{y}),$ (2)
@@
$\vec{p}(\boldsymbol{x}\mid\boldsymbol{y})=\frac{\vec{p}(\boldsymbol{y}\boldsymbol{x})}{\vec{p}(\boldsymbol{y})}$ (3)
where we additionally define $\vec{p}(\textsc{eos}\mid\boldsymbol{y})\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\frac{p(\boldsymbol{y})}{\vec{p}(\boldsymbol{y})}$.
@@
where $\boldsymbol{y}_{<n}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}y_{1}\cdots y_{n-1}$ and $\textsc{eos}\notin\Sigma$ is a distinguished end-of-string symbol. Let $\overline{\Sigma}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\Sigma\cup\{\textsc{eos}\}$. In Eq. 4, each $\vec{p}(\cdot\mid\boldsymbol{y}_{<n})$ can fruitfully be viewed as a distribution over $\overline{\Sigma}$. Despite the overloading of the notation, whether $\vec{p}(\cdot\mid\boldsymbol{y}_{<n})$ refers to a prefix probability, a function which takes an argument from $\Sigma^{*}$, or a distribution over $\overline{\Sigma}$ will always be clear from context. Throughout this paper, we use $\boldsymbol{Y}$ to represent the string-valued random variable sampled from $p$. When taking $M$ samples from $p$, we use $\boldsymbol{Y}^{(m)}$ to denote the $m^{\text{th}}$ sample.
@@
$\textsc{kl}(p\mid\mid q)\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\sum_{\boldsymbol{y}\in\Sigma^{*}}p(\boldsymbol{y})\log\frac{p(\boldsymbol{y})}{q(\boldsymbol{y})}=\operatorname*{\mathbb{E}}\left[f(\boldsymbol{Y})\right],$ (5)
where the function that we want to estimate is denoted with $f$, which in this case is $\log\frac{p(\boldsymbol{Y})}{q(\boldsymbol{Y})}$. Indeed, for the remainder of the paper, we define $f(\boldsymbol{Y})\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\log\frac{p(\boldsymbol{Y})}{q(\boldsymbol{Y})}$. Note that while the exact KL value is always non-negative, $f(\boldsymbol{Y})$ may be positive or negative. Importantly, the exact computation of Eq. 5 involves iterating over all strings in $\Sigma^{*}$.
@@ 2.2 Monte Carlo Estimation of KL
$\delta_{\textsc{mc}}$ $=\frac{1}{M}\sum_{m=1}^{M}\log\frac{p(\boldsymbol{Y}^{(m)})}{q(\boldsymbol{Y}^{(m)})}=\frac{1}{M}\sum_{m=1}^{M}f(\boldsymbol{Y}^{(m)}),$ (6a)
where $\boldsymbol{Y}^{(1)},\ldots,\boldsymbol{Y}^{(M)}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny i.i.d.}}}{{\sim}}}p$. Throughout the paper, we assume that the KL divergence is finite, i.e., $\textsc{kl}(p\mid\mid q)<\infty$. It is straightforward to show $\delta_{\textsc{mc}}$ is unbiased, i.e., $\operatorname*{\mathbb{E}}[\delta_{\textsc{mc}}]=\textsc{kl}(p\mid\mid q)$. The variance of this estimator is
$\mathop{\mathrm{Var}}\left[{\normalcolor\delta_{\textsc{mc}}}\right]$ $=\frac{1}{M^{2}}\mathop{\mathrm{Var}}\left[{\normalcolor\sum_{m=1}^{M}f(\boldsymbol{Y}^{(m)})}\right]=\frac{1}{M}\mathop{\mathrm{Var}}\left[{\normalcolor f(\boldsymbol{Y})}\right].$ (7a)
@@ 2.3 Control Variate Monte Carlo
A general technique to reduce the variance of any estimator is by introducing control variates (§8.2; Ross, 2002). In the context of estimating the KL divergence between two language models, control variate Monte Carlo was popularized by Schulman (2020) and is commonly implemented in many RLHF libraries (Havrilla et al., 2023; Hu et al., 2024; Sheng et al., 2024). Formally, a control variate is any function $g\colon\Sigma^{*}\to\mathbb{R}$ for which $G\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\operatorname*{\mathbb{E}}\left[g(\boldsymbol{Y})\right]$ can be efficiently computed. [^4] One could also consider control variates of the form $g\colon\Sigma^{*}\to\mathbb{R}^{d}$ for $d>1$ (Geffner and Domke, 2018). Armed with a control variate, we define the control variate Monte Carlo as []
@@
which directly translates to reducing the variance of the Monte Carlo estimator (compare Eq. 7a and Eq. 10c). The greater the correlation between $f$ and $g$, the more we reduce the variance. Note that the value of $\alpha$ may be estimated from a pilot sample when it cannot be computed analytically. [^5] Note that the control variate method can be straightforwardly extended to support multiple control variates.
@@ 3 Rao–Blackwellized Monte Carlo
$\delta_{\textsc{mc}}^{N}$ =$\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\frac{1}{M}\sum_{m=1}^{M}\log\frac{\vec{p}(\overline{\boldsymbol{Y}}_{\leq N}^{(m)})}{\vec{q}(\overline{\boldsymbol{Y}}_{\leq N}^{(m)})}$ (15a)
@@
$=\sum_{n=1}^{N}\underbrace{\frac{1}{M}\sum_{m=1}^{M}\log\frac{\vec{p}(\overline{Y}_{n}^{(m)}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})}{\vec{q}(\overline{Y}_{n}^{(m)}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})}}_{\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\delta_{\textsc{mc}}^{n}}$ (15d)
@@
$\delta_{\textsc{rb}}^{N}$ =$\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\sum_{n=1}^{N}\operatorname*{\mathbb{E}}\left[\delta_{\textsc{mc}}^{n}\mid T_{n}\right]=\sum_{n=1}^{N}\frac{1}{M}\sum_{m=1}^{M}\operatorname*{\mathbb{E}}_{\overline{Y}_{n}}\left[\log\frac{\vec{p}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})}{\vec{q}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})}\right],$ (17)
where the expectation in Eq. 17 is the exact KL divergence between the conditional distributions and is computed as: $\sum_{\overline{Y}_{n}\in\overline{\Sigma}}\vec{p}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})\log\frac{\vec{p}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})}{\vec{q}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})}$. The following theorem establishes that $\delta_{\textsc{rb}}^{N}$ is an unbiased estimator of $\delta_{\textsc{mc}}^{N}$, and furthermore, that $\delta_{\textsc{rb}}^{N}$ has lower or equal variance compared to $\delta_{\textsc{mc}}^{N}$.
@@
With regard to the estimator $\delta_{\textsc{rb}}^{N}$ defined in Eq. 17, the following two properties hold
@@
We then define the general case as $\delta_{\textsc{rb}}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\lim_{N\rightarrow\infty}\delta_{\textsc{rb}}^{N}$. The following theorem shows that $\delta_{\textsc{rb}}$ is an unbiased estimator of the KL divergence with variance less than or equal to $\delta_{\textsc{mc}}$.
@@
Notably, $\delta_{\textsc{rb}}$ is guaranteed to be non-negative. Each step-wise estimator computes the exact KL divergence between the two distributions, conditioned on the sampled context $\boldsymbol{y}_{<n}^{(m)}$. Since all terms in Eq. 17 are non-negative, $\delta_{\textsc{rb}}$ remains non-negative as well.
@@
At first glance, computing the Rao–Blackwellized estimator in Eq. 17 might seem more computationally expensive than the standard Monte Carlo estimator in Eq. 6a, since it requires evaluating the exact KL divergence at each token position. However, this added computation does not change the overall runtime complexity. To sample $M$ strings of at most length $N$, we perform a forward pass through the language model for $N$ tokens, followed by the final transformation and a softmax. Assuming that the model representations live in $\mathbb{R}^{d}$, the next token prediction step requires ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathcal{O}\!\left(MNd|\Sigma|\right)}$ operations, where $|\Sigma|$ is the size of the alphabet. Thus, the total time complexity of sampling $M$ strings is ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathcal{O}\!\left(MT_{N}+MNd|\Sigma|\right)}$, where $T_{N}$ denotes the time complexity of performing a forward pass for a single string of length $N$. For KL estimation, computing the MC estimator requires an additional ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathcal{O}\!\left(MN\right)}$ operations, while the Rao–Blackwellized estimator requires ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathcal{O}\!\left(MN|\Sigma|\right)}$. Therefore, despite the additional computation in the Rao–Blackwellized estimator, its overall runtime complexity is the same as the standard Monte Carlo estimator. In general, $N$ is itself a random variable. Thus, a more fine-grained, randomized analyze would yield the bound ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathcal{O}\!\left(MT_{\mathbb{E}[N]}+M\mathbb{E}[N]d|\Sigma|\right)}$ where $\mathbb{E}[N]$ is the expected string length under $p$.
@@ 4 Estimating the Gradient
$\nabla_{\boldsymbol{\theta}}\textsc{kl}$ $=\sum_{\boldsymbol{y}\in\Sigma^{*}}\nabla_{\boldsymbol{\theta}}p_{\boldsymbol{\theta}}(\boldsymbol{y})\log\frac{p_{\boldsymbol{\theta}}(\boldsymbol{y})}{q(\boldsymbol{y})}+\nabla_{\boldsymbol{\theta}}p_{\boldsymbol{\theta}}(\boldsymbol{y})=\operatorname*{\mathbb{E}}\left[\left(\log\frac{p_{\boldsymbol{\theta}}(\boldsymbol{Y})}{q(\boldsymbol{Y})}+1\right)\nabla_{\boldsymbol{\theta}}\log p_{\boldsymbol{\theta}}(\boldsymbol{Y})\right].$ (19)
@@
$\nabla_{\boldsymbol{\theta}}\textsc{kl}(p_{\boldsymbol{\theta}}\mid\mid q)=\sum_{\boldsymbol{y}\in\Sigma^{*}}\vec{p}_{\boldsymbol{\theta}}(\boldsymbol{y})\operatorname*{\mathbb{E}}_{\overline{Y}}\Big{[}\log\frac{\vec{p}_{\boldsymbol{\theta}}(\overline{Y}\mid\boldsymbol{y})}{\vec{q}(\overline{Y}\mid\boldsymbol{y})}\cdot\left(\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\overline{Y}\mid\boldsymbol{y})+\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\boldsymbol{y})\right)\Big{]}.$ (23)
@@
$\nabla\delta_{\textsc{rb}}=\frac{1}{M}\sum_{m=1}^{M}\lim_{N\rightarrow\infty}\sum_{n=1}^{N}\operatorname*{\mathbb{E}}_{\overline{Y}_{n}}\Big{[}\log\frac{\vec{p}_{\boldsymbol{\theta}}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})}{\vec{q}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})}\cdot\left(\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\overline{\boldsymbol{Y}}^{(m)}_{<n})+\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\overline{Y}_{n}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})\right)\Big{]}.$ (24)
@@
Let $p_{\boldsymbol{\theta}}$ and $q$ be two language models over $\Sigma$ and $\vec{p}_{\boldsymbol{\theta}}$ the prefix probability function of $p_{\boldsymbol{\theta}}$. Let $\nabla\delta_{\textsc{mc}}$ be the Monte Carlo estimator of the true gradient, denoted with $G\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\nabla_{\boldsymbol{\theta}}\textsc{kl}(p_{\boldsymbol{\theta}}\mid\mid q)$, as defined in Eq. 21 and $\nabla\delta_{\textsc{rb}}$ be the Rao–Blackwellized estimator of the gradient defined in Eq. 24. Assuming $\mathrm{H}(p_{\boldsymbol{\theta}},q)<\infty,\mathop{\mathrm{Var}}\left[{\normalcolor\nabla\delta_{\textsc{rb}}}\right]<\infty,\mathop{\mathrm{Var}}\left[{\normalcolor\nabla\delta_{\textsc{mc}}}\right]<\infty$, we have